OpenStudy (anonymous):

An architect needs to determine the slope between two points on a ski lift. The two points have been identified as (10, 35) and (150, 55), where x is the horizontal distance and y is the vertical distance from the bottom of the lift. Assuming the lift runs in a straight line, what is the slope of the line between the two points?

jimthompson5910 (jim_thompson5910):

Are you familiar with the slope formula?

OpenStudy (anonymous):

not so much.the teacher was trying to teach it to me today and i told her i did not understand and she just moved on and told me to come to saturday school -.-

jimthompson5910 (jim_thompson5910):

that's ok

jimthompson5910 (jim_thompson5910):

Start off by subtracting the y coordinates: 55-35 = 20 --------------------------------------- Now subtract the x coordinates (in the same order): 150-10 = 140 --------------------------------------- So the y difference is 20 The x difference is 140 The slope is then 20/140 = 1/7

jimthompson5910 (jim_thompson5910):

Basically the slope is the fraction of change in y ----------- change in x OR y difference ----------- x difference

OpenStudy (anonymous):

o_0 im confused

jimthompson5910 (jim_thompson5910):

where are you stuck

jimthompson5910 (jim_thompson5910):

point it out please

OpenStudy (anonymous):

starting at the part where you where talking about the differences of y and x

OpenStudy (anonymous):

ooooooo wait never mind i understand!!! thank you!!!

jimthompson5910 (jim_thompson5910):

You basically subtract, which results in a difference Alright, I'm glad it's starting to click